PKTOGF(1) PKTOGF(1) NAME pktogf-convertpackedfontfilestogenericfontfiles SYNOPSIS pktogf[-v]pkfilename[gffilename] DESCRIPTION Thepktogfprogramconvertsapackedfontfile(PK)toagenericfontfile(GF).Packedfontfiles aremuchsmallerthanthecorrespondingGFfiles,butsomeDVIreadersonlyunderstandGF files. Thepkfilenameonthecommandlinemustbecomplete.Becausetheresolutionispartofthe extension,itwouldnotmakesensetoaddonadefaultextensionasisdonewithTeXorDVIreadingsoftware.Thegffilenamedefaultstothesame(stripped)nameaspkfilename,andit isplacedinthecurrentworkingdirectorywiththe‘gf’suffixreplacing‘pk’.Forexample,the inputfileio.300pkwouldbecomeio.300gf. OPTIONS Unlessthe-vswitchisgiven,pktogfoperatessilently.With-v,theversionnumberoftheprogram andstatisticsabouttheunpackingareoutputtostdout. ENVIRONMENT pktogflooksforpkfilenamefirstinthecurrentdirectory.Ifitisnotpresentthere,itusesthe environmentvariablePKFONTS.Ifthatisnotset,itusesthevariableTEXPKS.Ifthatisnot set,itusesTEXFONTS.Ifthisisnotset,itusesthesystemdefaults: TEXFONTS:/usr/TeX/lib/tex/fonts//:., Seetex(1)forthedetailsofthesearching. SEEALSO gftopk(1),gftype(1),pktype(1) DonaldKnuthetal.,METAFONTware AUTHORS TomasRokickiwrotetheprogram.PierreMacKayadapteditforcompilationwithweb2c. 7Jan92 1